OpenStudy (anonymous):

i need help would greatly appreciate it 8x+4-2=-35

OpenStudy (anonymous):

8x+4-2=-35 (add like terms) 8x+2=-35 (get x by itself) -2 -2 8x=-37 (divide) 8 8 x=-4.6